A year ago Denise was 4 times as old as her brother Bryan. This year she is only 3 times as old as Bryan. How old were Denise and Bryan last year?
Answer by stanbon(75887)

A year ago Denise was 4 times as old as her brother Bryan. This year she is only 3 times as old as Bryan. How old were Denise and Bryan last year?
------------------------------------------------
Let Bryan's age NOW be "B".
Let Denise's age NOW be "D".
------------------------------
Equations:
(D-1) = 4(B-1)
D = 3B
--------------------
Substitute for D to solve for B:
(3B-1) = 4(B-1)
3B-1 = 4B-4
B = 3 (Bryan's age NOW)
---
Substitute to solve for "D":
D = 3B = 3*3 = 9 (Denise's age NOW)
-------
Bryan's age last year was B-1 = 2
Denise's age last year was D-1 = 8
=========================================================
Cheers,
Stan H.